The following information is from Bowin Inc. for a long-term construction project that is expected to be completed in January 2021. The construction project is for a building intended for the company’s own use. The capital expenditure on January 1, 2020, is for the purchase of land for the building site. No new construction loans were opened for the project in 2020. All debt was outstanding for the full year.

Capital Expenditures for 2020

Date Amount
Jan. 1, 2020 $ 12,000
Mar. 31, 2020 360,000
June 30, 2020 720,000
Nov. 30, 2020 360,000

Outstanding Debt in 2020

Debt Debt Amount Interest Rate
Note payable $400,000 8%
Note payable 320,000 8%
Bond payable 800,000 10%
Note payable 200,000 9%

Answer the following questions:
a. Compute interest to be capitalized and the interest to be expensed in 2020.
b. Prepare the entry to record the construction expenditures and interest for 2020.
c. Prepare the entry for depreciation in 2021 assuming that the project is completed on January 1, 2021. Assume that the building has a useful life of 30 years, and that the company uses the straight-line depreciation method.


PLEASE SHOW YOU CALCULATIONS  Note: Do not round until your final answers, then round to nearest whole number.
Answer #1

Sol :

a)

Calculation of Interest to be Capitalized & Interest to be Expensed.

1.Calculation of Borrowing Cost as per AS-23 Borrowing Cost based on Capitalisation rate used for Non-Specific Borrowings.

8% Note Payable = $400,000*8%*12/12 =$32000

8% Note Payable = $320,000*8%*12/12 =$25600

10% Bond Payable = $800,000*10%*12/12 =$80000

9% Note Payable = $200,000*9%*12/12 = $18000

Total Borrowing Cost =$155600

Capitalisation Rate = Total Borrowing Cost / Total Funds

= ($155600 / (400000+320000+800000+200000))*100

= ($155600 / $1720000)*100

= 9.04%

Borrwing cost to be Capitalized

Jan01,2020 $12000*9.04%*12/12 = $1084.80

Mar31,2020 $360000*9.04%*9/12 = $24408

June30,2020 $720000*9.04%*6/12 = $32544

Nov30,2020 $360000*9.04%*1/12 = $2712

Borrowing Cost to be Capitalised = $60749

Interest to Capitalised = $60750(Rounding off )

Interest to be Expensed ($155600-$60750) = $94850

b)

Journal Entry

Expenditure to be Capitalised = (12000+360000+720000+360000) = $1452000

Interest to be Capitalised = $60750

Building A/c................Dr 1512750

To Cash/Bank 1512750

(Being capital expenditure & interest been capitalised (1452000+60750))

c)

Depreciation = (Cost - Salvage Value)/ Useful Life

= (1512750-0) / 35

= 43221.42

Depreciation = 43220

Accumulated Depreciation...............Dr 43220

To Building Cost 43220

Notes :

Capitalisation of Interest is made for 12 Months (from Jan,2020 - 1st Jan ,2021)

Borrowing cost to be allocated based on Expenditure Incurred from Non- Specific Borrowings.